OpenStudy (anonymous):

Algebra/Probability Two distinct integers, x and y, are randomly chosen from the set {1, 2, 3, 4, 5, 6, 7, 8, 9 ,10}. What is the probability that xy-x-y is even?

OpenStudy (anonymous):

I think splitting this up into some cases will help. Case 1: Suppose x and y are even. Then \(xy-x-y\) is also even, since (even\(\times\)even=even) and subtracting evens from an even gives an even. Case 2: Suppose x and y are odd. Then \(xy-x-y\) is even, since (odd\(\times\)odd=even) and subtracting an odd twice gives an even. Case 3: Suppose x is even and y is odd. Then \(xy-x-y\) is even, since (odd\(\times\)even=odd). Subtracting an odd gives an even, then subtracting an even gives another even. Looks like \(xy-x-y\) is even no matter what you pick for \(x\) and \(y\).

OpenStudy (anonymous):

But the probability is not 1

OpenStudy (anonymous):

I tried it, and it told me it was wrong.

OpenStudy (anonymous):

Maybe the accepted answer is 100%?

hartnn (hartnn):

xy-x-y+1 should be odd (x-1)(y-1) should be odd

OpenStudy (anonymous):

I don't understand how that works into the problem

hartnn (hartnn):

both x and y must be odd, right ?

OpenStudy (anonymous):

@hartnn is saying that the difference between an even and an odd number (consecutive numbers) is 1. Add one to \(xy-x-y\) and you have \(xy-x-y+1\), which is odd. Which means \(xy-x-y\) is even. I'm not misinterpreting am I?

hartnn (hartnn):

thats correct. you are not misintrepreting

OpenStudy (anonymous):

Ah, I see

hartnn (hartnn):

so only one case, both x and y must be odd, the probability of which is ?

OpenStudy (anonymous):

@wgary, hartnn's right, I made a mistake with my second case. (odd\(\times\)odd=odd\), not even. Subtracting two odds gives another odd. Sorry

OpenStudy (anonymous):

Hmm, so does that mean the probability would be 2/3?

OpenStudy (anonymous):

Wait, no, because there are uneven probabilities of each case occurring, right?

OpenStudy (anonymous):

I think the probability would be\[\frac{\text{(number of pairs of odds)}}{\text{(number of possible pairs)}}\]

hartnn (hartnn):

how u got 2/3 ?

OpenStudy (anonymous):

Or equivalently \(\dfrac{5\times4}{10\times9}\)

hartnn (hartnn):

i think it this way : choosing one odd from 10 -----> 5/10 = 1/2 AND choosing another odd from remaining nine, 4/9 total = 1/2* 4/9 = .... ?

OpenStudy (anonymous):

And so I'm assuming that means the probability would be 2/9? (of getting an even number)

OpenStudy (mathmate):

There are 4 cases with equal probabilities, three of which are even. even-even = even odd-odd = odd even-odd = even odd-even = even can you now figure out the probability of even?

hartnn (hartnn):

well, i'd say don't assume, prove it like we did :)

OpenStudy (mathmate):

@SithsAndGiggles already did most of the work with the four cases plus your correction. All we have to prove is that the probability of getting P(even) from {1,2,3,4,5,6,7,8,9,10} equals 5/10.

OpenStudy (anonymous):

Just to make sure, (with SithsandGiggles method), it would be 5 odds, then 4 odds for the next choice, and then 10 (even or odd number) and 9 for (even or odd number)

OpenStudy (mathmate):

@wgary We are choosing from the same pool of 10 numbers for both x and y. So the probability of choosing an even is the same for each of x and y.

OpenStudy (anonymous):

They're not being replaced though. x and y are distinct.

OpenStudy (anonymous):

I thought they were distinct numbers, so they couldn't be the same

hartnn (hartnn):

yup, 'distinct' is the keyword, they can't be same

OpenStudy (mathmate):

Oh, sorry, didn't read properly the question, my bad!

OpenStudy (anonymous):

Also, I think another way to think of it would be 5c2/10c2

OpenStudy (mathmate):

In that case, there will be 8 cases, the four sithsAndGiggles cases multiplied by odd first and even first. Symmetry tells me that the results are not changed.

hartnn (hartnn):

C (5,2) = 10 C (10,2) = 45 yes, 5C2 /10C2 will work :)

OpenStudy (mathmate):

I guess I goofed again. The probability of odd-odd is (5/10)*(4/9)=20/90=2/9 So P(even)=1-P(odd)=7/9.

OpenStudy (anonymous):

@SithsAndGiggles As hartnn said, 5c2 = 10, and 10c2 =45

hartnn (hartnn):

P = favorable events/ total events = choosing 2 odds from 5 / choosing 2 numbers from 10 = 5c2/10c2 :)

OpenStudy (anonymous):

Right, I was thinking of C(5,1) and C(10,1)... Sorry about that

OpenStudy (anonymous):

Well, thanks to everyone ;)

hartnn (hartnn):

welcome ^_^